OpenStudy (anonymous):

integrate: dt/sqrt(1-16t^2)

OpenStudy (anonymous):

Your original equation\[\int\limits_{}^{}\frac{ 1 }{ \sqrt{1-16t^2} }dt\]We use "u" substitution to make this problem easier to do. Let:\[u = 4t\]and:\[du=4dt\]then look up the differential of sin^-1(u) and see if there's a resemblance.